Sunday 12/22/13

9:45a – 11:50p


  • align 750-647 G1 double stains



  • update menu options for feducial drift correction to allow finding the bin file of feducial data using the browser (oops, this was already the default for an empty input, should add a note of that).
  • fixed MultiChnDriftCorrect (this was all sorts of crazy)
  • switched storm buttons to using previously uploaded buttons with ‘Visible’ ‘on’ ‘off’ property (note ‘on’ ‘off’ not equal to true false).
  • updated file loading so that STORMrender forces you to list the files in the order taken when loading. This will make things easier for chromewarp and drift correction to know which transforms to apply to which mlist.
  • meanwhile latest drift, chromatic warp gives reasonable alignment: see post

Issues with feducial drift correction

  • The image-to-image accuracy is not so good. The smoothed positions look much better.
  • It would be better still to bin frames. Maybe I should write a new dax movie. We could even toss the old dax movie and save the ‘compressed’ version.



new approach to drift correction

  • downsampling STORM movies
  • Lets modify ReadDax to add ROI finder first.
  • I suppose this doesn’t help compression unless we are willing to wait for splitDax to cut off the original.
  • Still need to figure out how I want to integrate this into the work-flow.

Ph project

  • antibodies for Ajaz
  • antibody #74, anti-m, 2:2
  • antibody #?, anti-rb, 2.8:1.3
  • antibody #? recent. cy7 anti-m, 5:5, .17 mg/mL
  • antibody #? recent. cy7 anti-rb.

Probe Making

  • D12, F11, F12, G1, G2, G3, G5, G6,
  • F03, F04, F05, F06, G04
  • Ran PCR
  • Finished PCR cleanup. Moved samples back to PCR tubes for IVT reactions. Now at -20C, numbered 1-13.
  • Old gel, gel 1 (D12-G06) didn’t run too well. Should make fresh stuff tomorrow.
  • Gel 2 melted (need to change 100% of the buffer to run back-to-back gels!) This would have been F03-G04
This entry was posted in Summaries and tagged , . Bookmark the permalink.